OpenStudy (anonymous):

What is 5/8 divided by 15/16? A.5/12 B.3/4 C.3/2 D.2/3

OpenStudy (texaschic101):

when dividing fractions, flip what you are dividing by, then multiply. (5/8) / (15/16) = 5/8 * 16/15 =